# Port-Hamiltonian Neural ODE Networks on Lie Groups For Robot Dynamics Learning and Control

### v1

@@ VII Evaluation
(d) $\mathbf{D}_{\omega}({\boldsymbol{\mathfrak{q}}})/\beta$ versus $\varphi$.
@@
VII-A Pendulum
We consider a pendulum with the following dynamics:
$\ddot{\varphi}=-15\sin{\varphi}+3u-0.2\dot{\varphi},$ (92)
representing the pendulum orientation. We let $\boldsymbol{\omega}=[0,0,\dot{\varphi}]$ and remove position $\mathbf{p}$ and linear velocity $\mathbf{v}$ from the Hamiltonian dynamics in (IV-E), restricting the system to the $SO(3)$ manifold with generalized coordinates $\mathbf{{\boldsymbol{\mathfrak{q}}}}=[\mathbf{r}_{1}^{\top}\quad\mathbf{r}_{2}^{\top}\quad\mathbf{r}_{3}^{\top}]^{\top}$.
As noted in [23, 27], since the generalized momenta $\mathbf{{\boldsymbol{\mathfrak{p}}}}$ are not available in the dataset, the dynamics of $\mathbf{{\boldsymbol{\mathfrak{q}}}}$ in (92) do not change if $\mathbf{{\boldsymbol{\mathfrak{p}}}}$ is scaled by a factor $\beta>0$. This is also true in our formulation as scaling $\mathbf{{\boldsymbol{\mathfrak{p}}}}$ leaves the dynamics of $\mathbf{{\boldsymbol{\mathfrak{q}}}}$ in (IV-E) unchanged. To emphasize this scale-invariance, let $\mathbf{M}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\beta\mathbf{M}(\mathbf{{\boldsymbol{\mathfrak{q}}}})$, ${\cal V}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\beta{\cal V}(\mathbf{{\boldsymbol{\mathfrak{q}}}})$, $\mathbf{D}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})=\beta\mathbf{D}(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})$ $\mathbf{B}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\beta\mathbf{B}(\mathbf{{\boldsymbol{\mathfrak{q}}}})$, and:
guaranteeing that the equations of motions (IV-E) still hold.
Fig. 3 shows the training and testing behavior of our $SO(3)$ Hamiltonian ODE network. Fig. 2(a) and 2(c) show that the $\left[\mathbf{M}(\mathbf{{\boldsymbol{\mathfrak{q}}}})^{-1}\right]_{3,3}$ entry of the mass inverse and the $\left[\mathbf{B}(\mathbf{{\boldsymbol{\mathfrak{q}}}})\right]_{3}$ entry of the input matrix with scaling factor $\beta=0.75$ are close to their correct values of $3$ and $1$, respectively, while the other entries are close to zero. Fig. 2(b) indicates a constant gap between the learned and the ground-truth potential energy, which can be explained by the relativity of potential energy.
@@ VII-A Pendulum
The controlled angle $\varphi$ and angular velocity $\dot{\varphi}$ as well as the control inputs $\mathbf{u}$ with gains $\mathbf{K}_{\mathbf{R}}=2\mathbf{I}$ and $\mathbf{K}_{\mathbf{d}}=\mathbf{I}$ are shown over time in Fig. 2(e) and 2(f). We can see that the controller was able to smoothly drive the pendulum from $\phi=0$ to $\phi=\pi$, relying only on the learned dynamics.
VII-B Crazyflie Quadrotor
The quadrotor was controlled from a random starting point to $18$ different desired poses using a PID controller [72], providing $18$ $2.5$-second trajectories. The trajectories were used to generate a dataset $\mathcal{D}=\{t_{0:N}^{(i)},\mathbf{{\boldsymbol{\mathfrak{q}}}}_{0:N}^{(i)},\boldsymbol{\zeta}_{0:N}^{(i)},\mathbf{u}^{(i)})\}_{i=1}^{D}$ with $N=5$ and $D=1080$. The $SE(3)$ port-Hamiltonian ODE network was trained, as described in Sec. V-D, for $500$ iterations without a nominal model, i.e., $\mathbf{M}^{-1}_{\mathbf{v}0}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\bf 0$, $\mathbf{M}^{-1}_{\boldsymbol{\omega}0}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\bf 0$, $\mathbf{D}_{\mathbf{v}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $\mathbf{D}_{\boldsymbol{\omega}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $V_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$ and $\mathbf{B}_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$. We do not consider energy dissipation such as drag effect in the PyBullet simulator, and we omit the dissipation matrix $\mathbf{D}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$ in the model design.
@@ VII-B Crazyflie Quadrotor
Figure 4: Evaluation of $SE(3)$ port-Hamiltonian neural ODE network on a Crazyflie quadrotor in the PyBullet simulator [72].
@@
Figure 8: Comparison of different neural network architecture to learn quadrotor dynamics: 1) black-box, i.e., the dynamics function $\mathbf{f}$ is represented by a multilayer perceptron network; 2) unstructured Hamiltonian, i.e., the Hamiltonian function is represented by a multilayer perceptron network instead of the sum of kinetic and potential energy as shown in Eq. (27); 3) structured Hamiltonian, i.e., the Hamiltonian function has the form of Eq. (27). Initialized at a random pose and twist, the learned quadrotor dynamics are rolled out, showing that our approach with structured Hamiltonian preserves $SO(3)$ constraints and the conservation of energy better than the other models, and provides better state predictions.
@@
VII-C Comparison to Unstructured Neural ODE Models
@@
In this section, we show the benefits of our neural ODE architecture by comparing 1) our structured Hamiltonian model, 2) a black-box model, i.e., the approximated dynamics $\mathbf{f}$ is represented by a multilayer perceptron network, and 3) an unstructured Hamiltonian model, i.e., the Hamiltonian function is represented by a multilayer perceptron network instead of using the structure in Eq. (27), in terms of training convergence rates, guarantees of energy conservation principle, and Lie group constraint satisfaction. To verify energy conservation, we rolled out the learned dynamics and calculated the Hamiltonian via (27) along the predicted trajectories for (1) the black-box model using ground-truth mass and potential energy with the predicted states; (2) the unstructured Hamiltonian model using the output of the multilayer perceptron Hamiltonian network; and (3) the structured Hamiltonian model using the learned mass and potential energy networks. We check the $SO(3)$ constraints by verifying that two quantities $|\det{\mathbf{R}}-1|$ and $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$ remain small along the predicted trajectories.
We first use a pendulum as described in Sec. VII-A without energy dissipation. The models are trained for $5000$ iterations from $512$ 0.2-second state-control trajectories and rolled out for a significantly longer horizon of $50$ seconds. Fig. 7 plots the training loss, the phase portraits, the SO(3) constraints and the total energy (Hamiltonian) of the learned models for a pendulum system. As the Hamiltonian structure is imposed in the neural ODE network architecture, our model is able to converge faster with lower loss (Fig. 6(a)) and preserves the phase portraits for state predictions (Fig. 6(b)). Fig. 6(c) shows that the $SO(3)$ constraints are satisfied by our structured and unstructured Hamiltonian models as their values of $|\det{\mathbf{R}}-1|$ and $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$ remain small along a $50$-second trajectory rollout initialized at $\phi=\pi/2$. The constant Hamiltonian in Fig. 6(d) of our structured Hamiltonian verifies that our model obeys the energy conservation law with no control input and no energy dissipation. The Hamiltonian of the black-box model increases along the trajectory while that of the unstructured Hamiltonian model fluctuates and slightly decreases over time.
We also tested the models using the simulated Crazyflie quadtoror with the same dataset ${\cal D}$ of $18$ trajectories as described in Sec. VII-B. The $SE(3)$ port-Hamiltonian ODE network was trained, as described in Sec. V-D, for $500$ iterations. Our structured Hamiltonian model converges faster with significantly lower loss as seen in Fig. 7(a). We verified that the predicted orientation trajectories from our learned models satisfy the $SO(3)$ constraints. Fig. 7(b) shows two near-zero quantities $|\det{\mathbf{R}}-1|$ and $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$, obtained by rolling out our learned dynamics for $5$ seconds, while the learned black-box model violates the constraints after a very short time. Fig. 7(c) shows a constant total energy along the predicted trajectory from our structured Hamiltonian model without control input and dissipation networks, verifying that the learned model obeys the law of energy conservation. Fig. 7(d) shows that our structured Hamiltonian model provides better trajectory predictions compared to the other methods.
@@
Refer to caption (d) Tracking at $t=0$s Refer to caption (e) Tracking at $t=3$s Refer to caption (f) Tracking at $t=9$s Refer to caption (g) Tracking at $t=12$s
@@
VII-D Real Quadrotor Experiments
VII-D1 Learning quadrotor dynamics after upgrade
We consider a scenario in which the quadrotor is upgraded with a new frame and a new onboard computer, leading to changes in the robot dynamics that we aim to learn from data. The mass and inertia matrix of a previous model (Fig. 8(a)) with Raspberry Pi computer and an F450 frame were used as the nominal mass $\mathbf{M}_{\mathbf{v}0}=1.3\mathbf{I}$ and inertia matrix $\mathbf{M}_{\boldsymbol{\omega}0}=\operatorname{diag}([0.12,0.12,0.2])$ for the upgraded quadrotor (Fig. 8(b)). The other nominal matrices were set to zero: $\mathbf{D}_{\mathbf{v}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $\mathbf{D}_{\boldsymbol{\omega}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $V_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$ and $\mathbf{B}_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$. We modified the PX4 firmware [73] to expose the normalized thrust and torque being sent to the motors. The firmware’s normalization of thrust and torque is unknown and, in fact, is learned from data via the input gain matrix $\mathbf{B}({\boldsymbol{\mathfrak{q}}})$. We collected $12$ state-control trajectories by flying the quadrotor from a starting pose to $12$ different poses using a PID controller provided by the PX4 flight controller [73]. The trajectories were used to generate a dataset $\mathcal{D}=\{t_{0:N}^{(i)},\mathbf{{\boldsymbol{\mathfrak{q}}}}_{0:N}^{(i)},\boldsymbol{\zeta}_{0:N}^{(i)},\mathbf{u}^{(i)})\}_{i=1}^{D}$ with $N=1$ and $D=~{}10000$. We trained our model as described in Sec. V-D for $5000$ steps.
The trained model was used with the control policy in Sec. VII-B to track different trajectories: a verticle circle, a vertical lemniscate, and a 3D piecewise-linear trajectories. Fig. 10 and 11 show that we achieve better tracking performance using our learned dynamics model and energy-based control compared to the nominal model and the geometric controller in [70]. The tracking errors of our controller with a learned model improve by $2-4$ times compared to those of geometric control based on the nominal model, as shown in Table I.
TABLE I: Position tracking errors using nominal and learned quadrotor models with and without payload.
@@ VII-D1 Learning quadrotor dynamics after upgrade
VII-D2 Learning quadrotor dynamics with extra payload
In this section, we demonstrate that after our dynamics model is trained, if there is a change in the quadrotor dynamics, e.g., an extra payload is added, we are able to update the dynamics quickly starting from the previously trained model. We attached a coffee can to the quadrotor frame (Fig. 8(c)) to change the mass and inertia matrix of the robot. We, then, collected a new dataset by driving the quadrotor to $12$ different poses, and trained our dynamics model for only $100$ steps, initialized with the trained model in Sec. VII-D1.
In the presence of the coffee can payload, the tracking performance of the controller with the previously learned model degrades as shown in Fig. 12 and 13. Meanwhile, after a quick model update, the robot is able to track desired trajectories accurately again. Table I shows that our updated model improves the tracking errors by 1.51.51.51.54444 times compared to the previously learned model.

### v2

@@ VII Evaluation
VII-A Pendulum
In this section, we verify our port-Hamiltonian dynamics learning and control approach on the $SO(3)$ manifold. We consider a pendulum with the following dynamics:
$\ddot{\varphi}=-15\sin{\varphi}+3u-0.2\dot{\varphi},$ (68)
representing the pendulum orientation. We let $\boldsymbol{\omega}=[0,0,\dot{\varphi}]$ and remove position $\mathbf{p}$ and linear velocity $\mathbf{v}$ from the Hamiltonian dynamics in (30), restricting the system to the $SO(3)$ manifold with generalized coordinates $\mathbf{{\boldsymbol{\mathfrak{q}}}}=[\mathbf{r}_{1}^{\top}\quad\mathbf{r}_{2}^{\top}\quad\mathbf{r}_{3}^{\top}]^{\top}$.
As noted in [27, 32], since the generalized momenta $\mathbf{{\boldsymbol{\mathfrak{p}}}}$ are not available in the dataset, the dynamics of $\mathbf{{\boldsymbol{\mathfrak{q}}}}$ in (68) do not change if $\mathbf{{\boldsymbol{\mathfrak{p}}}}$ is scaled by a factor $\beta>0$. This is also true in our formulation as scaling $\mathbf{{\boldsymbol{\mathfrak{p}}}}$ leaves the dynamics of $\mathbf{{\boldsymbol{\mathfrak{q}}}}$ in (30) unchanged. To emphasize this scale-invariance, let $\mathbf{M}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\beta\mathbf{M}(\mathbf{{\boldsymbol{\mathfrak{q}}}})$, ${\cal V}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\beta{\cal V}(\mathbf{{\boldsymbol{\mathfrak{q}}}})$, $\mathbf{D}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})=\beta\mathbf{D}(\mathbf{{\boldsymbol{\mathfrak{q}}}},\mathbf{{\boldsymbol{\mathfrak{p}}}})$ $\mathbf{B}_{\beta}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\beta\mathbf{B}(\mathbf{{\boldsymbol{\mathfrak{q}}}})$, and:
guaranteeing that the equations of motions (30) still hold.
@@ VII-A Pendulum
(d) $\beta\mathbf{D}_{\omega}({\boldsymbol{\mathfrak{q}}})$ versus $\varphi$.
@@
(a) Loss (log scale)
(b) $\mathbf{M}_{\mathbf{v}}^{-1}({\boldsymbol{\mathfrak{q}}})$ and $\mathbf{M}_{\boldsymbol{\omega}}^{-1}({\boldsymbol{\mathfrak{q}}})$
(c) ${\cal V}({\boldsymbol{\mathfrak{q}}})$
(d) $\mathbf{B}_{\mathbf{v}}({\boldsymbol{\mathfrak{q}}})$
(e) SO(3) constraints.
(f) Total energy.
Figure 4: Evaluation of $SE(2)$ port-Hamiltonian neural ODE network on a simulated omnidirectional ground vehicle with scale factor $\beta=7.1$.
Fig. 3 shows the training and testing behavior of our $SO(3)$ Hamiltonian ODE network. Fig. 3(a) and 3(c) show that the $\left[\mathbf{M}(\mathbf{{\boldsymbol{\mathfrak{q}}}})^{-1}\right]_{3,3}$ entry of the mass inverse and the $\left[\mathbf{B}(\mathbf{{\boldsymbol{\mathfrak{q}}}})\right]_{3}$ entry of the input matrix with scaling factor $\beta=1.33$ are close to their correct values of $3$ and $1$, respectively, while the other entries are close to zero. Fig. 3(b) indicates a constant gap between the learned and the ground-truth potential energy, which can be explained by the relativity of potential energy.
@@
The controlled angle $\varphi$ and angular velocity $\dot{\varphi}$ as well as the control inputs $\mathbf{u}$ with gains $\mathbf{K}_{\mathbf{R}}=2\mathbf{I}$ and $\mathbf{K}_{\mathbf{d}}=\mathbf{I}$ are shown over time in Fig. 3(e) and 3(f). We can see that the controller was able to smoothly drive the pendulum from $\phi=0$ to $\phi=\pi$, relying only on the learned dynamics.
VII-B Omnidirectional Ground Vehicle
In this section, we verify our port-Hamiltonian dynamics learning and control approach on a simulated omnidirectional ground vehicle, whose states evolve on the $SE(2)$ manifold. The ground-truth dynamics of the vehicle can be obtained from (30) by keeping only the components $x$ and $y$ in the position $\mathbf{p}$ and the yaw angle of the rotation matrix, leading to an $SO(2)$ rotation matrix:
$\mathbf{R}=\begin{bmatrix}\cos{\varphi}&-\sin{\varphi}\\ \sin{\varphi}&\cos{\varphi}\end{bmatrix},$ (72)
where $\varphi$ is the vehicle’s yaw angle. The vehicle moves on a flat ground with potential energy ${\cal V}({\boldsymbol{\mathfrak{q}}})=c$, where $c$ is a constant, and has ground-truth mass $\mathbf{M}_{\mathbf{v}}({\boldsymbol{\mathfrak{q}}})=\mathbf{I}\in\mathbb{S}_{\succ 0}^{2\times 2}$ and inertia $\mathbf{M}_{\boldsymbol{\omega}}({\boldsymbol{\mathfrak{q}}})=0.05\in\mathbb{R}_{>0}$. It is fully-actuated with control input $\mathbf{u}=[f_{x},f_{y},\tau_{\varphi}]$, where $f_{x}$ and $f_{y}$ are forces along the $x$ and $y$ axes of the body frame and $\tau_{\varphi}$ is the yaw torque, generated by the motors.
(a) Tracking performance with a lemniscate trajectory.
(b) Tracking lemniscate (left) and piecewise-linear (right) trajectories.
Figure 5: Trajectories (blue) of omnidirectional ground vehicle tracking desired trajectories (orange) with our learned $SE(2)$ port-Hamiltonian dynamics and IDA-PBC control design.
@@ VII-B Omnidirectional Ground Vehicle
Figure 6: $SE(3)$ port-Hamiltonian neural ODE network on a Crazyflie quadrotor in the PyBullet simulator [81].
To collect training data, the vehicle was controlled from a random initial point to $9$ different desired positions and yaw angles using a PID controller, providing $9$ one-second trajectories. The trajectories were used to generate a dataset $\mathcal{D}=\{t_{0:N}^{(i)},\mathbf{{\boldsymbol{\mathfrak{q}}}}_{0:N}^{(i)},\boldsymbol{\zeta}_{0:N}^{(i)},\mathbf{u}^{(i)})\}_{i=1}^{D}$ with $N=5$ and $D=432$. The $SE(2)$ port-Hamiltonian ODE network was formulated, as described in Sec. V-D, by ignoring the $z$ component of the position $\mathbf{p}$, and the pitch and roll components of the rotation $\mathbf{R}$. The model was trained for $8000$ iterations without a nominal model, i.e., $\mathbf{M}^{-1}_{\mathbf{v}0}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\bf 0$, $\mathbf{M}^{-1}_{\boldsymbol{\omega}0}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\bf 0$, $\mathbf{D}_{\mathbf{v}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $\mathbf{D}_{\boldsymbol{\omega}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $V_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$ and $\mathbf{B}_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$. We did not consider energy dissipation such as friction in the simulation, and we omitted the dissipation matrix $\mathbf{D}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$ in the model.
Fig. 4 shows the training results for the $SE(2)$ Hamiltonian ODE network. Fig. 4(b) and 4(d) show that the mass inverse and the input gain matrix with scaling factor $\beta=7.1$ are close to their correct values: $\mathbf{M}_{\mathbf{v}}(\mathbf{{\boldsymbol{\mathfrak{q}}}})^{-1}\approx\mathbf{I},\mathbf{M}_{\boldsymbol{\omega}}(\mathbf{{\boldsymbol{\mathfrak{q}}}})^{-1}\approx 20$ and $\mathbf{B}(\mathbf{{\boldsymbol{\mathfrak{q}}}})\approx\mathbf{I}$. Fig. 4(c) shows a constant learned potential energy as expected.
We verified our energy-based control design in Sec. VI by controlling the ground robot to track horizontal lemniscate and piecewise-linear trajectories. Fig. 5(b) demonstrates that the omnidirectional ground vehicle controlled by our energy-based controller achieves successful trajectory tracking. The control gains were chosen as: $\mathbf{K}_{\mathbf{p}}=0.72\mathbf{I}$, $\mathbf{K}_{\mathbf{v}}=0.8\mathbf{I},\mathbf{K}_{\mathbf{R}}=9.1\mathbf{I}$, $\mathbf{K}_{\boldsymbol{\omega}}=3.6$. Fig. 5(a) plots the tracking errors in the position, yaw angles, linear velocity, and angular velocity.
VII-C Crazyflie Quadrotor
The quadrotor was controlled from a random starting point to $18$ different desired poses using a PID controller [81], providing $18$ $2.5$-second trajectories. The trajectories were used to generate a dataset $\mathcal{D}=\{t_{0:N}^{(i)},\mathbf{{\boldsymbol{\mathfrak{q}}}}_{0:N}^{(i)},\boldsymbol{\zeta}_{0:N}^{(i)},\mathbf{u}^{(i)})\}_{i=1}^{D}$ with $N=5$ and $D=1080$. The $SE(3)$ port-Hamiltonian ODE network was trained, as described in Sec. V-D, for $500$ iterations without a nominal model, i.e., $\mathbf{M}^{-1}_{\mathbf{v}0}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\bf 0$, $\mathbf{M}^{-1}_{\boldsymbol{\omega}0}(\mathbf{{\boldsymbol{\mathfrak{q}}}})=\bf 0$, $\mathbf{D}_{\mathbf{v}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $\mathbf{D}_{\boldsymbol{\omega}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $V_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$ and $\mathbf{B}_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$. We did not consider energy dissipation such as drag effect in the PyBullet simulator, and we omitted the dissipation matrix $\mathbf{D}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$ in the model design.
@@ VII-C Crazyflie Quadrotor
Figure 10: Comparison of different neural network architectures to learn quadrotor dynamics: 1) black-box, i.e., the dynamics function $\mathbf{f}$ is represented by a multilayer perceptron network; 2) unstructured Hamiltonian, i.e., the Hamiltonian function is represented by a multilayer perceptron network instead of the sum of kinetic and potential energy as shown in Eq. (27); 3) structured Hamiltonian, i.e., the Hamiltonian function has the form of Eq. (27). Initialized at a random pose and twist, the learned quadrotor dynamics are rolled out, showing that our approach with structured Hamiltonian preserves $SO(3)$ constraints and the conservation of energy better than the other models, and provides better state predictions.
@@
VII-D Comparison to Unstructured Neural ODE Models
@@
(d) Tracking at $t=0$s
(e) Tracking at $t=3$s
(f) Tracking at $t=9$s
(g) Tracking at $t=12$s
@@
In this section, we show the benefits of our neural ODE architecture by comparing 1) our structured Hamiltonian model, 2) a black-box model, i.e., the approximated dynamics $\mathbf{f}$ is represented by a multilayer perceptron network, and 3) an unstructured Hamiltonian model, i.e., the Hamiltonian function is represented by a multilayer perceptron network instead of using the structure in Eq. (27), in terms of training convergence rates, satisfaction of energy conservation principle, and Lie group constraints. To verify energy conservation, we rolled out the learned dynamics and calculated the Hamiltonian via (27) along the predicted trajectories for (1) the black-box model using ground-truth mass and potential energy with the predicted states; (2) the unstructured Hamiltonian model using the output of the multilayer perceptron Hamiltonian network; and (3) the structured Hamiltonian model using the learned mass and potential energy networks. We check the $SO(3)$ constraints by verifying that two quantities $|\det{\mathbf{R}}-1|$ and $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$ remain small along the predicted trajectories.
We first use a pendulum as described in Sec. VII-A without energy dissipation. The models are trained for $5000$ iterations from $512$ 0.2-second state-control trajectories and rolled out for a significantly longer horizon of $50$ seconds. Fig. 9 plots the training loss, the phase portraits, the SO(3) constraints and the total energy (Hamiltonian) of the learned models for a pendulum system. As the Hamiltonian structure is imposed in the neural ODE network architecture, our model is able to converge faster with lower loss (Fig. 9(a) and Table I), preserves the phase portraits for state predictions (Fig. 9(b)), and achieves the lowest angle prediction error in in Table I. Fig. 9(c) and Table I show that the $SO(3)$ constraints are satisfied by our structured and unstructured Hamiltonian models as their values of $|\det{\mathbf{R}}-1|$ and $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$ remain small along a $50$-second trajectory rollout initialized at $\phi=\pi/2$. The constant Hamiltonian in Fig. 9(d) of our structured Hamiltonian with lowest standard derivation value in Table I verifies that our model obeys the energy conservation law with high precision, given no control input and no energy dissipation. The Hamiltonian of the black-box model increases along the trajectory while that of the unstructured Hamiltonian model fluctuates and slightly decreases over time.
We also tested the models using the simulated Crazyflie quadtoror with the same dataset ${\cal D}$ of $18$ trajectories as described in Sec. VII-C. The $SE(3)$ port-Hamiltonian ODE network was trained, as described in Sec. V-D, for $500$ iterations. Our structured Hamiltonian model converges faster with significantly lower loss as seen in Fig. 10(a) and in Table I. We verified that the predicted orientation trajectories from our learned models satisfy the $SO(3)$ constraints. Fig. 10(b) and Table I show two near-zero quantities $|\det{\mathbf{R}}-1|$ and $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$, obtained by rolling out our learned dynamics for $5$ seconds, while the learned black-box model significantly violates the constraints after a very short time. Fig. 10(c) shows a constant total energy along the predicted trajectory, i.e., lowest standard derivation value in Table I, from our structured Hamiltonian model without control input and dissipation networks, verifying that the learned model obeys the law of energy conservation. Fig. 10(d) and the prediction error in Table I show that our structured Hamiltonian model provides better trajectory predictions compared to the other methods.
TABLE I: Comparison of different neural network architectures for pendulum and quadrotor dynamics learning.
Metrics Platform Black-box Unstructured Structured Hamiltonian Hamiltonian Training loss Pendulum $0.037$ $6.3\times 10^{-5}$ $\mathbf{2.3\times 10^{-7}}$ $\|det(\mathbf{R})-1\|$ (avg.) Pendulum $1888752$ $1.4\times 10^{-3}$ $\mathbf{1.4\times 10^{-3}}$ $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$ (avg.) Pendulum $267971.6$ $2.1\times 10^{-3}$ $\mathbf{2.1\times 10^{-3}}$ Total energy (std.) Pendulum $13557.1$ $0.163$ $\mathbf{0.003}$ Prediction error (avg.) Pendulum $1.02~{}(rad)$ $0.08~{}(rad)$ $\mathbf{0.008~{}(rad)}$ Training loss Quadrotor $2.2\times 10^{-3}$ $6.4\times 10^{-4}$ $\mathbf{3.92\times 10^{-6}}$ $\|det(\mathbf{R})-1\|$ (avg.) Pendulum $3741$ $2.9\times 10^{-6}$ $\mathbf{2.6\times 10^{-7}}$ $\|\mathbf{R}\mathbf{R}^{\top}-\mathbf{I}\|$ (avg.) Quadrotor $29336.7$ $7.6\times 10^{-6}$ $\mathbf{1.3\times 10^{-6}}$ Total energy (std.) Quadrotor $18.1$ $0.074$ $\mathbf{1.64\times 10^{-6}}$ Prediction error (avg.) Quadrotor $0.49~{}(m)$ $0.46~{}(m)$ $\mathbf{0.02~{}(m)}$
TABLE II: Position tracking errors using nominal and learned quadrotor models with and without payload.
@@
VII-E Real Quadrotor Experiments
VII-E1 Learning quadrotor dynamics after upgrade
We consider a scenario in which the quadrotor is upgraded with a new frame and a new onboard computer, leading to changes in the robot dynamics that we aim to learn from data. The nominal model was obtained from a computer-aided design (CAD) model of another much lighter-weight Raspberry Pi quadrotor with a Raspberry Pi computer and an F450 frame (Fig. 11(a)), which is less accurate and far from the unknown ground-truth model of our upgraded quadrotors. Specifically, the nominal mass and inertia matrix are $\mathbf{M}_{\mathbf{v}0}=1.3\mathbf{I}$ and $\mathbf{M}_{\boldsymbol{\omega}0}=\operatorname{diag}([0.12,0.12,0.2])$, respectively, for the upgraded quadrotor in Fig. 11(b). The other nominal matrices were set to zero: $\mathbf{D}_{\mathbf{v}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $\mathbf{D}_{\boldsymbol{\omega}0}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, $V_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$ and $\mathbf{B}_{0}({\boldsymbol{\mathfrak{q}}})=\bf 0$. We modified the PX4 firmware [82] to expose the normalized thrust and torque being sent to the motors. The firmware’s normalization of thrust and torque is unknown and, in fact, is learned from data via the input gain matrix $\mathbf{B}({\boldsymbol{\mathfrak{q}}})$. We collected $12$ state-control trajectories by flying the quadrotor from a starting pose to $12$ different poses using a PID controller provided by the PX4 flight controller [82]. The trajectories were used to generate a dataset $\mathcal{D}=\{t_{0:N}^{(i)},\mathbf{{\boldsymbol{\mathfrak{q}}}}_{0:N}^{(i)},\boldsymbol{\zeta}_{0:N}^{(i)},\mathbf{u}^{(i)})\}_{i=1}^{D}$ with $N=1$ and $D=~{}10000$. We trained our model as described in Sec. V-D for $5000$ steps.
The trained model was used with the control policy in Sec. VII-C to track different trajectories: a verticle circle, a vertical lemniscate, and a 3D piecewise-linear trajectories. Fig. 12 and 13 show that we achieve better tracking performance using our learned dynamics model and energy-based control compared to the nominal model and the geometric controller in [79]. The tracking errors of our controller with a learned model improve by $2-4$ times compared to those of geometric control based on the nominal model, as shown in Table II.
VII-E2 Learning quadrotor dynamics with extra payload
In this section, we demonstrate that after our dynamics model is trained, if there is a change in the quadrotor dynamics, e.g., an extra payload is added, we are able to update the dynamics quickly starting from the previously trained model. We attached a coffee can to the quadrotor frame (Fig. 11(c)) to change the mass and inertia matrix of the robot. We, then, collected a new dataset by driving the quadrotor to $12$ different poses, and trained our dynamics model for only $100$ steps, initialized with the trained model in Sec. VII-E1.
In the presence of the coffee can payload, the tracking performance of the controller with the previously learned model degrades as shown in Fig. 14 and 15. Meanwhile, after a quick model update, the robot is able to track desired trajectories accurately again. Table II shows that our updated model improves the tracking errors by 1.51.51.51.54444 times compared to the previously learned model.